# Decentralised Finance and Automated Market Making: Predictable Loss and Optimal Liquidity Provision

### v2

@@ 1 Introduction
Existing research characterises the losses of LPs, but does not offer tools for strategic liquidity provision. In this paper, we study strategic liquidity provision in CPMs with CL. We derive the continuous-time dynamics of the wealth of strategic LPs which consists of the position they hold in the pool (position value), fee income, and costs from repositioning their liquidity position. The width of the range where the assets are deposited affects the value of the LP’s position in the pool; specifically, we show that the predictable loss (PL) incurred by LPs increases as the width of the liquidity range decreases. PL measures the unhedgeable losses of LPs stemming from the depreciation of their holdings in the pool and from the opportunity costs from locking their assets in the pool; see Cartea et al. (2023a). PL is similar to loss-versus-rebalancing (LVR) in Milionis et al. (2022) which describes the unhedgeable losses of LPs in traditional CFMs due to losses to arbitrageurs. [^1] Milionis et al. (2022) introduced LVR in August 2022 as a measure that quantifies the unhedgeable losses of LPs in CFMMs to arbitrageurs. Contemporaneously, Cartea et al. (2023a) introduced PL in November 2022 as a component in the wealth of strategic LPs that quantifies predictable losses due to the convexity of the trading function and due to opportunity costs in CFMs and in CL markets. Also, we show that fee income is subject to a tradeoff between the width of the LP’s liquidity range and the volatility of the marginal rate in the pool. More precisely, CL increases fee revenue when the rate is in the range of the LP, but also increases concentration risk. Concentration risk refers to the risk that the LP faces when her position is concentrated in narrow ranges; the LP stops collecting fees when the rate exits the range of her position.
@@
Early works on AMMs are in Chiu and Koeppl (2019), Angeris et al. (2021), Lipton and Treccani (2021). Some works in the literature study strategic liquidity provision in CFMs and CPMs with CL. Heimbach et al. (2022) discuss the tradeoff between risks and returns that LPs face in Uniswap v3, Cartea et al. (2023a) study the predictable losses of LPs in a continuous-time setup, Milionis et al. (2023) study the impact of fees on the profits of arbitrageurs in CFMs, Fukasawa et al. (2023) study the hedging of the impermanent losses of LPs, and Lı et al. (2023) study the economics of liquidity provision. Closest to our work are the models in Fan et al. (2021) and Fan et al. (2022) which focus on fee revenue and use approximation techniques to obtain dynamic strategies. Finally, there is a growing literature on AMM design for fair competition between LPs and LTs. Goyal et al. (2023) study an AMM with dynamic trading functions that incorporate beliefs of LPs, Lommers et al. (2023) study AMMs where the LP’s strategy adjusts dynamically to market information, and Cartea et al. (2023b) generalise CFMs and propose AMM designs where LPs express their beliefs and risk preferences.
@@ 2.1 Constant function markets
$f\left(q^{X}+(1-\tau)\,x,q^{Y}-y\right)=\kappa^{2}\quad\implies\quad\tilde{Z}(y)=\frac{\varphi_{\kappa}\left(q^{Y}-y\right)-\varphi_{\kappa}(y)}{\left(1-\tau\right)\,y}\,.$ (1)
@@
Liquidity provision operations in CPMs do not impact the marginal rate, so when an LP deposits the quantities $x$ and $y$ of assets $X$ and $Y$, the condition $Z=q^{X}/q^{Y}=(q^{X}+x)/(q^{Y}+y)$ must be satisfied; see Cartea et al. (2022a, b).
@@ 2.2 Concentrated liquidity markets
In CPMs with CL, the proportional fee $\tau$ charged by the pool to LTs is distributed among LPs. More precisely, if an LP’s liquidity position with depth $\tilde{\kappa}$ is in a tick range where the total depth of liquidity is $\kappa$, then for every liquidity taking trade that pays an amount $p$ of fees, the LP with liquidity $\tilde{\kappa}$ earns the amount
@@ 3.1 Position value
where $\delta^{\ell}\in\left(-\infty,2\right]$, $\delta^{u}\in\left[-\infty,2\right)$, and $\delta^{\ell}\,\delta^{u}/2<\delta^{\ell}+\delta^{u}$ because $0\leq Z^{\ell}<Z^{u}<\infty.$ We define $\delta^{\ell}$ and $\delta^{u}$ in (6) in terms of $\sqrt{Z}$ to simplify and linearise the CL constant product formulae; see Cartea et al. (2023a) for more details.
@@
We assume that the marginal rate process $\left(Z_{t}\right)_{t\in[0,T]}$ follows the dynamics (5). Cartea et al. (2023a) show that the holdings in assets $X$ and $Y$ in the pool for an LP who follows an arbitrary strategy $\left(Z_{t}^{\ell},Z_{t}^{u}\right)$ are given by
@@
where the predictable and negative component $\text{PL}_{t}=-\frac{\sigma^{2}}{2}\,\int_{0}^{t}\frac{\tilde{x}_{s}}{\delta_{s}}\,\mathrm{d}s$ is the PL of the LP’s position which scales with the volatility of the marginal rate. PL is related to LVR (see Milionis et al. (2022)) which decomposes the loss of LPs in traditional CFMs into a hedgeable market risk and an unhedgeable component due to profits made by arbitrageurs. The dynamics in (3.1) show that a larger position spread $\delta$ reduces PL and the overall risk of the LP’s position in the pool; see Cartea et al. (2023a) for more details.
@@ 3.2.1 Fee income: pool fee rate
Table 1: First row: correlation of the returns of the rate $Z$ and the fee rate $\pi$, i.e., $\left(Z_{t+\Delta t}-Z_{t}\right)/Z_{t}$ and $\left(Z_{t+\Delta t}-Z_{t}\right)/Z_{t}$ for $\Delta t=1$ minute, five minutes, one hour, and one day, using data of the ETH/USDC pool between 5 May 2021 and 18 August 2022. Second row: two-tailed p-value of the t-test. Final row: $R$-squared of regression of the pool fee rate’s returns against the marginal exchange rate’s returns.
@@ 4.3 Discussion: profitability, PL, and concentration risk
When $\delta^{\ell}=\delta^{u}=2,$ the LP provides liquidity in the maximum range $(0\,,+\infty),$ so the depth of her liquidity position $\tilde{\kappa}$ is minimal, the PL is minimal, and the LP’s position is equivalent to providing liquidity in CPMs without CL; see Cartea et al. (2023a) for more details. In that case, the dynamics of PL in (3.1) are
@@ 5.1 Methodology
In this section, we use Uniswap v3 data between 1 January and 18 August 2022 (see data description in A) to study the performance of the strategy of Section 4. We consider execution costs and discuss how gas fees and liquidity taking activity in the pool affect the performance of the strategy. [^12] In practice, LPs pay gas fees when using the Ethereum network to deposit liquidity, withdraw liquidity, and adjust their holdings. Gas is paid in Ether, the native currency of the Ethereum network, and measures the computational effort of the LP operation; see Cartea et al. (2022a).
@@
To compute the LP’s performance as a result of changes in the value of her holdings in the pool (position value), and as a result of fee income, we use out-of-sample data $[t_{i},t_{i+1}].$ We use equation (3.1) to determine the one-minute out-of-sample changes in the position value. For fee income, we use LT transactions in the pool at rates included in the range $\left(Z_{t}^{\ell},Z_{t}^{u}\right]$ and equation (4). The income from fees accumulates in a separate account in units of $X$ with zero risk-free rate. [^13] In practice, fees accumulate in both assets $X$ and $Y.$ At the end of the out-of-sample window, the LP withdraws her liquidity and collects the accumulated fees, and we repeat the in-sample estimation and out-of-sample liquidity provision described above. Thus, at times $t_{i},$ where $i\in\{1,\dots,N-1\},$ the LP consecutively withdraws and deposits liquidity in different ranges. Between two consecutive operations (i.e., reposition liquidity provision), the LP may need to take liquidity in the pool to adjust her holdings in asset $X$ and $Y.$ In that case, we use results in Cartea et al. (2022a) to compute execution costs. [^14] Cartea et al. (2022a) show that execution costs in the pool are a closed-form function of the rate $Z,$ the pool depth $\kappa,$ and the transaction size. In particular, we consider execution costs when the LP trades asset $Y$ in the pool to adjust her holdings between two consecutive operations. More precisely, we consider that for every quantity $y$ of asset $Y$ bought or sold in the pool, a transaction cost $y\,Z_{t}^{3/2}/\kappa$ is incurred. We assume that the LP’s taking activity does not impact the dynamics of the pool. Finally, we obtain 331,858 individual LP operations from 1 January to 18 August 2022.
@@ 6 Discussion: modelling assumptions
Our model assumes continuous repositioning of the LP’s position. However, when interacting with blockchains, updates occur at the block validation frequency. For instance, the Ethereum network’s blocks are validated every 13 seconds, on average. Moreover, within each block, the transactions form a (random) queue that determines the priority with which they are executed. Our model can be extended to include delays inherent to blockchains and we refer the reader to the work in Cartea and Sánchez-Betancourt (2021) and Cartea et al. (2021).
@@
Our model assumes that gas fees paid by the LP to interact with the blockchain are flat and constant. In practice, gas fees are stochastic and depend on exogenous factors such as the price of electricity and network congestion. These could be included in our model by considering a stochastic gas fee that is observed by the LP.

### v3

# Forthcoming in SIAM Journal on Financial Mathematics

@@ 1 Introduction
Existing research characterises the losses of LPs, but does not offer tools for strategic liquidity provision. In this paper, we study strategic liquidity provision in CPMs with CL. We derive the continuous-time dynamics of the wealth of strategic LPs which consists of the position they hold in the pool (position value), fee income, and costs from repositioning their liquidity position. The width of the range where the assets are deposited affects the value of the LP’s position in the pool; specifically, we show that the predictable loss (PL) incurred by LPs increases as the width of the liquidity range decreases. PL measures the unhedgeable losses of LPs stemming from the depreciation of their holdings in the pool and from the opportunity costs from locking their assets in the pool; see Cartea et al. (2023b). PL is similar to loss-versus-rebalancing (LVR) in Milionis et al. (2022) which describes the losses in traditional CFMs when LPs hedge their price exposure in an exogenous market. [^1] Milionis et al. (2022) introduced LVR in August 2022 as a measure that quantifies the unhedgeable losses of LPs in CFMMs when prices form in a centralised venue. Contemporaneously, Cartea et al. (2023b) introduced PL in November 2022 as a component in the wealth of strategic LPs that quantifies losses due to the convexity of the trading function and due to opportunity costs in CFMs and in CL markets. Also, we show that fee income is subject to a tradeoff between the width of the LP’s liquidity range and the volatility of the marginal rate in the pool. More precisely, CL increases fee revenue when the rate is in the range of the LP, but also increases concentration risk. Concentration risk refers to the risk that the LP faces when her position is concentrated in narrow ranges; the LP stops collecting fees when the rate exits the range of her position.
@@
Early works on AMMs are in Chiu and Koeppl (2019), Angeris et al. (2021), Lipton and Treccani (2021), Capponi and Jia (2021), Engel and Herlihy (2021a, b), Angeris et al. (2022); see also the recent review Biais et al. (2023). Capponi et al. (2023a) study price formation in AMMs. Some works in the literature study strategic liquidity provision in CFMs and CPMs with CL. Heimbach et al. (2022) discuss the tradeoff between risks and returns that LPs face in Uniswap v3, Cartea et al. (2023b) study the predictable losses of LPs in a continuous-time setup, Milionis et al. (2023) study the impact of fees on the profits of arbitrageurs in CFMs, Fukasawa et al. (2023) study the hedging of the impermanent losses of LPs, and Lı et al. (2023) study the economics of liquidity provision. Closest to our work are the models in Fan et al. (2021) and Fan et al. (2022) which focus on fee revenue and use approximation techniques to obtain dynamic strategies. Finally, there is a growing literature on AMM design for fair competition between LPs and LTs. Goyal et al. (2023) study an AMM with dynamic trading functions that incorporate beliefs of LPs, Lommers et al. (2023) study AMMs where the LP’s strategy adjusts dynamically to market information, and Cartea et al. (2023c) generalise CFMs and propose AMM designs where LPs express their beliefs and risk preferences; see also Bergault et al. (2024) and He et al. (2024).
@@ 2.1 Constant function markets
$f\left(q^{X}+(1-\tau)\,x,q^{Y}-y\right)=\kappa^{2}\quad\implies\quad\tilde{Z}(y)=\frac{\varphi_{\kappa}\left(q^{Y}-y\right)-\varphi_{\kappa}(q^{Y})}{\left(1-\tau\right)\,y}\,.$ (1)
@@
Liquidity provision operations in CPMs do not impact the marginal rate, so when an LP deposits the quantities $x$ and $y$ of assets $X$ and $Y$, the condition $Z=q^{X}/q^{Y}=(q^{X}+x)/(q^{Y}+y)$ must be satisfied; see Cartea et al. (2022, 2023b).
@@ 2.2 Concentrated liquidity markets
In CPMs with CL, the proportional fee $\tau$ charged by the pool to LTs is distributed among LPs. More precisely, if an LP’s liquidity position with depth $\tilde{\kappa}$ is in a tick range where the total depth of liquidity is $\kappa$, then for every liquidity taking trade that pays an amount $p$ of fees, [^4] The proportional fee $\tau$ represents a fixed percentage applied to the trade size to calculate the fee $p$ paid by an LT. Thus, the amount $p$ is expressed in terms of the reference asset $X$. the LP with liquidity $\tilde{\kappa}$ earns the amount
@@ 3.1 Position value
where $\delta^{\ell}\in\left(-\infty,2\right]$, $\delta^{u}\in\left[-\infty,2\right)$, and $\delta^{\ell}\,\delta^{u}/2<\delta^{\ell}+\delta^{u}$ because $0\leq Z^{\ell}<Z^{u}<\infty.$ We define $\delta^{\ell}$ and $\delta^{u}$ in (6) in terms of $\sqrt{Z}$ to simplify and linearise the CL constant product formulae; see Cartea et al. (2023b) for more details.
@@
We assume that the marginal rate process $\left(Z_{t}\right)_{t\in[0,T]}$ follows the dynamics (5). Cartea et al. (2023b) show that the holdings in assets $X$ and $Y$ in the pool for an LP who follows an arbitrary strategy $\left(Z_{t}^{\ell},Z_{t}^{u}\right)$ are given by
@@
where the predictable and negative component $\text{PL}_{t}=-\frac{\sigma^{2}}{2}\,\int_{0}^{t}\frac{\tilde{x}_{s}}{\delta_{s}}\,\mathrm{d}s$ is the PL of the LP’s position which scales with the volatility of the marginal rate. PL is related to LVR (see Milionis et al. (2022)) which decomposes the loss of LPs in traditional CFMs into a hedgeable market risk and an unhedgeable component due to profits made by arbitrageurs. The dynamics in (3.1) show that a larger position spread $\delta$ reduces PL and the overall risk of the LP’s position in the pool; see Cartea et al. (2023b) for more details.
@@ 3.2.1 Fee income: pool fee rate
In contrast to the proportional fee $\tau$, which represents the fixed percentage applied to the trade sizes of LTs to compute the fees paid to the pool, the pool fee rate $\pi$ denotes the instantaneous profitability of the pool computed as the percentage of total fees paid by LTs relative to the size of the pool. Thus, the pool fee rate $\pi$ depends on the intensity of the liquidity taking flow, the size of the pool, and the proportional fee $\tau$. Below, we also consider the fee revenue $p$ of one LP who maximises her wealth by choosing an optimal spread for her liquidity position.
@@
Table 1: First row: correlation of the returns of the rate $Z$ and the fee rate $\pi$, i.e., $\left(Z_{t+\Delta t}-Z_{t}\right)/Z_{t}$ and $\left(\pi_{t+\Delta t}-\pi_{t}\right)/\pi_{t}$ for $\Delta t=1$ minute, five minutes, one hour, and one day, using data of the ETH/USDC pool between 5 May 2021 and 18 August 2022. Second row: two-tailed p-value of the t-test. Final row: $R$-squared of regression of the pool fee rate’s returns against the marginal exchange rate’s returns.
@@ 4.3 Discussion: profitability, PL, and concentration risk
When $\delta^{\ell}=\delta^{u}=2,$ the LP provides liquidity in the maximum range $(0\,,+\infty),$ so the depth of her liquidity position $\tilde{\kappa}$ is minimal, the PL is minimal, and the LP’s position is equivalent to providing liquidity in CPMs without CL; see Cartea et al. (2023b) for more details. In that case, the dynamics of PL in (3.1) are
@@ 5.1 Methodology
In this section, we use Uniswap v3 data between 1 January and 18 August 2022 (see data description in A) to study the performance of the strategy of Section 4. We consider execution costs and discuss how gas fees and liquidity taking activity in the pool affect the performance of the strategy. [^13] In practice, LPs pay gas fees when using the Ethereum network to deposit liquidity, withdraw liquidity, and adjust their holdings. Gas is paid in Ether, the native currency of the Ethereum network, and measures the computational effort of the LP operation; see Cartea et al. (2022, 2023a).
@@
To compute the LP’s performance as a result of changes in the value of her holdings in the pool (position value), and as a result of fee income, we use out-of-sample data $[t_{i},t_{i+1}].$ We use equation (3.1) to determine the one-minute out-of-sample changes in the position value. For fee income, we use LT transactions in the pool at rates included in the range $\left(Z_{t}^{\ell},Z_{t}^{u}\right]$ and equation (4). The income from fees accumulates in a separate account in units of $X$ with zero risk-free rate. [^14] In practice, fees accumulate in both assets $X$ and $Y.$ At the end of the out-of-sample window, the LP withdraws her liquidity and collects the accumulated fees, and we repeat the in-sample estimation and out-of-sample liquidity provision described above. Thus, at times $t_{i},$ where $i\in\{1,\dots,N-1\},$ the LP consecutively withdraws and deposits liquidity in different ranges. Between two consecutive operations (i.e., reposition liquidity provision), the LP may need to take liquidity in the pool to adjust her holdings in asset $X$ and $Y.$ In that case, we use results in Cartea et al. (2022) to compute execution costs. [^15] Cartea et al. (2022) show that execution costs in the pool are a closed-form function of the rate $Z,$ the pool depth $\kappa,$ and the transaction size. In particular, we consider execution costs when the LP trades asset $Y$ in the pool to adjust her holdings between two consecutive operations. More precisely, we consider that for every quantity $y$ of asset $Y$ bought or sold in the pool, a transaction cost $y\,Z_{t}^{3/2}/\kappa$ is incurred. We assume that the LP’s taking activity does not impact the dynamics of the pool. Finally, we obtain 331,858 individual LP operations from 1 January to 18 August 2022.
@@ 6 Discussion: modelling assumptions
Our model assumes continuous repositioning of the LP’s position. However, when interacting with blockchains, updates occur at the block validation frequency. For instance, the Ethereum network’s blocks are validated every 13 seconds, on average. Moreover, within each block, the transactions form a (random) queue that determines the priority with which they are executed. Finally, price formation in Ethereum blockchains leads to sandwich attacks; see Capponi et al. (2023b) for more details. Our model can be extended to include delays inherent to blockchains and we refer the reader to the work in Cartea and Sánchez-Betancourt (2021) and Cartea et al. (2021).
@@
Our model assumes that gas fees paid by the LP to interact with the blockchain are flat and constant. In practice, gas fees are stochastic and depend on exogenous factors such as the price of electricity and network congestion. These could be included in our model by considering a stochastic gas fee that is observed by the LP. Also, our model is optimal when the spread of liquidity position is continuously adjusted to account for changes in the stochastic drift and profitability of liquidity taking according to (33). Thus, profitable liquidity provision using our strategy requires a large initial wealth to overcome gas fees from continuous trading. However, we expect the strategy to remain profitable in discrete time when the stochastic drift $\mu$ and the stochastic profitability $\pi$ remain stable, so the LP only rebalances her liquidity position when either the drift $\mu$ or the pool fee rate $\pi$ undergo large changes.